# RegMean++: Enhancing Effectiveness and Generalization of Regression Mean for Model Merging

### v1

@@ A.1 Datasets
Task descriptions and statistics of datasets used for the standard $8$-task image classification benchmark [^1] https://huggingface.co/collections/tanganke/the-eight-image-classification-tasks-6644ce0376c0a469f6928507 are described as follows:
@@
  • SUN397 (Xiao et al. 2016) contains more than $100,000$ images of $397$ categories for benchmarking scene understanding. The number of images varies across categories, but there are at least $100$ images each.
@@
  • Stanford Cars (Cars) (Krause et al. 2013) has $16,185$ images in total of $196$ types of cars and evenly split for training and testing sets.
@@
  • RESISC45 (Cheng, Han, and Lu 2017) is developed for remote sensing image scene classification. This dataset covers $45$ scene classes with $700$ images of size $256\times 256$ for each.
@@
  • EuroSAT (Helber et al. 2019) is used for land use and land cover classification using Sentinel-2 satellite images of size $64\times 64$, consisting of $27,000$ images covering $10$ classes.
@@
  • SVHN (Netzer et al. 2011) is a street view house number classification benchmark, containing more than $600,000$ RGB images of $10$ printed digits in size $32\times 32$ cropped from house number plates.
@@
  • GTSRB (Stallkamp et al. 2011) is a German traffic sign recognition benchmark consisting of over $50,000$ images of $43$ classes of traffic signs in varying light and background conditions.
@@
  • MNIST (LeCun et al. 1998), a well-known classical dataset for hand-written digit classification with $60,000$ training and $10,000$ testing images of size $28\times 28$ in $10$ classes of numbers.
@@
  • DTD (Cimpoi et al. 2014) is a collection of $5,640$ images across $47$ categories of textures in the wild, annotated with human-centric attributes.
Additional datasets.
@@
Besides the standard $8$-task scenario, we follow the previous work and further extend our experimental scenario to $20$ tasks [^2] https://huggingface.co/collections/tanganke/image-classification-datasets-662abda7d75efe6b0e6b43da. The new $12$ tasks are listed as follows:
@@
  • Flowers102 (Nilsback and Zisserman 2008) contains $102$ flower categories that are popular in the United Kingdom, with $1,020$ training and $6,149$ testing images. The images have varying poses and light conditions.
@@
  • PCAM (PatchCamelyon) (Veeling et al. 2018) consists of more than 300300M color images in size of $96\times 96$ pixels extracted from histopathologic scans of lymph node sections. Each of them is annotated with a binary class indicating the presence of metastatic tissue.
@@
  • FER2013 (Goodfellow et al. 2013) is developed for facial expression recognition. The images are grayscale and have a size of $48\times 48$ pixels, describing seven different kinds of emotions. The training and testing split consists of $28,709$ and $7,178$ samples, respectively.
@@
  • OxfordIIITPet (Parkhi et al. 2012) is a $37$-category pet dataset with roughly 200 images for each category, and is equally divided for both training and testing splits. The images vary in scale, pose, and lighting conditions.
@@
  • STL10 (Coates, Ng, and Lee 2011) is primarily built for unsupervised image recognition tasks covering $10$ classes. Hence, the number of labeled images is quite small: $500$ training and $800$ testing images for each class. All of them are in $96\times 96$ pixel resolution.
@@
  • CIFAR100 (Krizhevsky et al. 2009) consists of color images categorized in $100$ general classes, each class contains $600$ images, and each image is in size $32\times 32$. There are $50,000$ training images and $10,000$ testing images.
  • CIFAR10 (Krizhevsky et al. 2009) is similar to CIFAR100, except it has $10$ classes.
@@
  • Food101 (Bossard, Guillaumin, and Van Gool 2014) contains of $101$ food categories, with $101,000$ images. For each class, $750$ images are for training and $250$ are for testing. Only the testing images are manually reviewed. The training images contain noise mostly from intense colors, and sometimes are mislabelled.
@@
  • FashionMNIST (Xiao, Rasul, and Vollgraf 2017) is designed as a drop-in replacement benchmark for the original MNIST, thereby inheriting the same structure as MNIST.
@@
  • EMNIST (Cohen et al. 2017) is an extended version of MNIST. EMNIST contains images of both characters and digits. We choose to use only the EMNIST Letters split, which contains around $145,000$ images evenly distributed in $26$ classes of the alphabet letters.
@@
  • KMNIST (Clanuwat et al. 2018), yet another version of MNIST, represents $10$ Japanese Hiragana characters.
@@
  • RenderedSST2 (Socher et al. 2013b; Radford et al. 2019) is used for evaluating the models’ capability on optical character recognition. The images are rendered from sentences in the Stanford Sentiment Treebank v2 (Socher et al. 2013a), with black texts on a white background in $448\times 448$ resolution. Each image is labeled as positive or negative based on the mood expressed in the text, and the number of images for both classes is nearly balanced. There are $6,920$ training and $1,821$ testing images.
A.2 Models
@@
We employ off-the-shelf fine-tuned checkpoints from the previous work (Tang et al. 2024a), covering three architectures of pre-trained CLIP model (Radford et al. 2021): ViT-B/32333https://huggingface.co/collections/tanganke/clip-vit-b-32-on-the-eight-image-classication-tasks-6644d0c476c0a469f693cf91, ViT-B/16444https://huggingface.co/collections/tanganke/clip-vit-b-16-on-the-eight-image-classification-tasks-66cd54d8332ce5c7468ab5f8, and ViT-L/14555https://huggingface.co/collections/tanganke/clip-vit-l-14-on-the-eight-image-classification-tasks-6644d2b014331c746683de63. The indicators $32$, $16$, and $14$ mean the size of patches in pixels that an input image is divided into. We only merge the vision encoding part of these architectures, while the text encoding part is kept unchanged. The number of parameters for the vision encoding part of these architectures is 87.587.5M, 85.885.8M, and 303303M, respectively.
A.3 Merging Methods
@@
    • Model Soups (Wortsman et al. 2022) is the most straightforward approach that simply takes the average of candidate models’ parameters to produce a merge model.
    • Task Arithmetic (Ilharco et al. 2022) introduces a concept called “task vector”, which is the difference between the fine-tuned model’s parameters and the pre-trained parameters. A multi-tasking task vector is defined as the sum of those task vectors and is scaled by a coefficient before being added back to the pre-trained model’s parameters to produce a merge model.
    • TIES-Merging (Yadav et al. 2023) proposes to trim the small values of task vectors, then resolve sign conflicts before adding back to the pre-trained parameters.
    • TSV-M (Gargiulo et al. 2025) compresses the task vectors using singular value decomposition (SVD) to reduce the interference between task vectors at the layer level before merging.
    • DOGE TA (Wei et al. 2025) is a variant of Task Arithmetic, where DOGE, an iterative algorithm minimizing the gap between the merge model and the candidates while retaining the shared knowledge, is integrated.
    • Iso-C and Iso-CTS (Marczak et al. 2025). The former applied SVD on the merge task vector to identify the directions amplified by multiple tasks, i.e., common subspace. The latter further incorporates task-specific subspaces for retaining unique task features.
@@
    • Fisher Merging (Matena and Raffel 2022) produces the merge models by taking the weighted average of candidate models, with the weighting factors determined by the Fisher information matrices.
    • RegMean (Jin et al. 2022) proposes a closed-form solution for merging multiple linear layers, then applies this idea to the transformer models.
@@
    • Layer-wise AdaMerging (Yang et al. 2024b) adaptively learns the merging coefficients introduced by Task Arithmetic in the layer-wise or task-wise manner by using unsupervised entropy minimization on unlabeled test datasets.
    • DOGE AM (Wei et al. 2025) is another variant of AdaMerging, where DOGE is integrated.
(a) Clean
(b) Motion Blur
(c) Impulse Noise
(d) Gaussian Noise
(e) Pixelate
(f) Spatter
(g) Contrast
(h) JPEG Compression
Figure 7: A visualization of a clean image and its corrupted versions with seven types of common noises (Hendrycks and Dietterich 2019).
Appendix B Experimental Details
B.1 Normalized Accuracy Metric
@@
B.2 Hyperparameters
@@
Both require two hyperparameters: the number of samples per task and the scaling factor $\alpha$. As illustrated in main text Figure 5 (left) and Figure 6, increasing the number of samples for each task gives a relatively small improvement on performance across architectures. We choose to use $256$ samples for calculating the task’s inner-product matrices with a batch size of $32$ as default. For the scaling factor $\alpha$, Table 6 shows that a higher $\alpha$ delivers better performance. However, when $\alpha=1.0$, i.e., no scaling applied, the degradation happens and the overall accuracy is almost zeroed out. This phenomenon is consistent with the insight from Jin et al. (2022). Therefore, $\alpha=0.95$ is the optimal value.
@@
  • For Task Arithmetic (Ilharco et al. 2022), the merging coefficient $\lambda=0.3$ is used.
  • For TIES-Merging (Yadav et al. 2023), top-$20\%$ highest-magnitude parameters are retained for each task vector, then these trimmed task vectors are merged with the merging coefficient $\lambda=0.3$.
  • For TSV-M (Gargiulo et al. 2025), the task scaling factor $\alpha=1.0$ is used.
  • For DOGE TA (Wei et al. 2025), the modification vector $\Delta$ is optimized on $400$ iterations with a learning rate 1e41e-4 via Adam optimizer  (Kingma and Ba 2014), the global magnitude of merging coefficient $\eta=0.07$, the shared subspace basis size is set as the rank of the shared subspace divided by $6$, and top-$30\%$ highest-magnitude parameters are retained for each task vector.
  • For Iso-C (Marczak et al. 2025), the task scaling factor $\alpha$ is set to $1.30$, $1.40$, and $1.50$ for ViT-B/32, ViT-B/16, and ViT-L/14, respectively.
  • For Iso-CTS (Marczak et al. 2025), the task scaling factor $\alpha$ is set to $1.50$, $1.60$, and $1.90$ for ViT-B/32, ViT-B/16, and ViT-L/14, respectively; the size of the common subspace is set as its rank multiplied by $0.8$.
  • For Fisher Merging (Matena and Raffel 2022), the number of samples per task is $256$ and the batch size is $32$ for calculating the Fisher information matrix.
  • For Layer-wise AdaMerging (Yang et al. 2024b), the Adam optimizer is used with a learning rate of 1e31e-3 for updating the merging coefficients on $1,000$ iterations with the batch size of $16$.
  • For DOGE AM (Wei et al. 2025), its hyperparameters are the same as DOGE TA and Layer-wise AdaMerging.
B.3 Details for Experimental Settings
Orders of tasks for sustainability evaluation.

### v2

@@ B.1 Datasets
Task descriptions and statistics of datasets used for the standard $8$-task image classification benchmark are described below. These datasets are publicly available https://huggingface.co/collections/tanganke/the-eight-image-classification-tasks.
@@
  • SUN397 (xiao2016sun) contains more than $100,000$ images of $397$ categories for benchmarking scene understanding. The number of images varies across categories, but there are at least $100$ images each.
@@
  • Stanford Cars (Cars) (krause20133d) has $16,185$ images in total of $196$ types of cars and evenly split for training and testing sets.
@@
  • RESISC45 (cheng2017remote) is developed for remote sensing image scene classification. This dataset covers $45$ scene classes with $700$ images of size $256\times 256$ for each.
@@
  • EuroSAT (helber2019eurosat) is used for land use and land cover classification using Sentinel-2 satellite images of size $64\times 64$, consisting of $27,000$ images covering $10$ classes.
@@
  • SVHN (netzer2011reading) is a street view house number classification benchmark, containing more than $600,000$ RGB images of $10$ printed digits in size $32\times 32$ cropped from house number plates.
@@
  • GTSRB (stallkamp2011german) is a German traffic sign recognition benchmark consisting of over $50,000$ images of $43$ classes of traffic signs in varying light and background conditions.
@@
  • MNIST (lecun1998gradient), a well-known classical dataset for hand-written digit classification with $60,000$ training and $10,000$ testing images of size $28\times 28$ in $10$ classes of numbers.
@@
  • DTD (cimpoi2014describing) is a collection of $5,640$ images across $47$ categories of textures in the wild, annotated with human-centric attributes.
Additional vision classification datasets.
@@
Besides the standard $8$-task scenario, we follow the previous work and further extend our experimental scenario to $20$ tasks. The new $12$ tasks are listed below. These datasets are publicly available at https://huggingface.co/collections/tanganke/image-classification-datasets.
@@
  • Flowers102 (nilsback2008automated) contains $102$ flower categories that are popular in the United Kingdom, with $1,020$ training and $6,149$ testing images. The images have varying poses and light conditions.
@@
  • PCAM (PatchCamelyon) (veeling2018rotation) consists of more than 300300M color images in size of $96\times 96$ pixels extracted from histopathologic scans of lymph node sections. Each of them is annotated with a binary class indicating the presence of metastatic tissue.
@@
  • FER2013 (goodfellow2013challenges) is developed for facial expression recognition. The images are grayscale and have a size of $48\times 48$ pixels, describing seven different kinds of emotions. The training and testing split consists of $28,709$ and $7,178$ samples, respectively.
@@
  • OxfordIIITPet (parkhi2012cats) is a $37$-category pet dataset with roughly 200 images for each category, and is equally divided for both training and testing splits. The images vary in scale, pose, and lighting conditions.
@@
  • STL10 (coates2011analysis) is primarily built for unsupervised image recognition tasks covering $10$ classes. Hence, the number of labeled images is quite small: $500$ training and $800$ testing images for each class. All of them are in $96\times 96$ pixel resolution.
@@
  • CIFAR100 (krizhevsky2009learning) consists of color images categorized in $100$ general classes, each class contains $600$ images, and each image is in size $32\times 32$. There are $50,000$ training images and $10,000$ testing images.
  • CIFAR10 (krizhevsky2009learning) is similar to CIFAR100, except it has $10$ classes.
@@
  • Food101 (bossard2014food) contains of $101$ food categories, with $101,000$ images. For each class, $750$ images are for training and $250$ are for testing. Only the testing images are manually reviewed. The training images contain noise mostly from intense colors, and sometimes are mislabelled.
@@
  • FashionMNIST (xiao2017fashion) is designed as a drop-in replacement benchmark for the original MNIST, thereby inheriting the same structure as MNIST.
@@
  • EMNIST (cohen2017emnist) is an extended version of MNIST. EMNIST contains images of both characters and digits. We choose to use only the EMNIST Letters split, which contains around $145,000$ images evenly distributed in $26$ classes of the alphabet letters.
@@
  • KMNIST (clanuwat2018deep), yet another version of MNIST, represents $10$ Japanese Hiragana characters.
@@
  • RenderedSST2 (socher2013recursive; radford2019language) is used for evaluating the models’ capability on optical character recognition. The images are rendered from sentences in the Stanford Sentiment Treebank v2 (socher-etal-2013-recursive), with black texts on a white background in $448\times 448$ resolution. Each image is labeled as positive or negative based on the mood expressed in the text, and the number of images for both classes is nearly balanced. There are $6,920$ training and $1,821$ testing images.
Language generation datasets.
We provide a detailed description of the $11$ datasets used for language generation evaluation as follows.
  • IFEval (zhou2023instruction) is a straightforward and easy-to-reproduce benchmark on instruction-following evaluation. It contains $541$ “verifiable instructions” such as “write in more than $400$ words” and “mention the keyword of AI at least $3$ times”. The dataset is publicly available at https://huggingface.co/datasets/google/IFEval.
  • GSM8K (cobbe2021training) stands for Grade School Math 88K, which contains $8,792$ high quality grade school math problems created by human writers. These problems take between $2$ and $8$ steps to solve, where the solutions primarily involve performing a sequence of basic arithmetic operations ($\times\div$). There are $1,319$ test problems. The dataset is publicly available at https://huggingface.co/datasets/openai/gsm8k.
  • Multilingual MMLU, Multilingual ARC, and Multilingual Hellaswag (lai-etal-2023-okapi) are the ChatGPT-translated versions from English of three corresponding datasets, i.e, MMLU (hendryckstest2021), ARC (clark2018think), and Hellaswag (zellers2019hellaswag). Although there are $26$ languages, following (he2025mergebench), we evaluate the merge models on French, Spanish, German, and Russian. All of these datasets are organized as multiple-choice question-answering tasks, which focus on different types of knowledge. MMLU assesses the model’s multi-task accuracy on a wide range of world knowledge and problem-solving ability. ARC challenges models on reasoning tasks, which comprise natural, grade-school science questions. Hellaswag provides commonsense natural language inference questions that are trivial for humans, but difficult for state-of-the-art models. These translated datasets are publicly provided as follows: Multilingual MMLU at https://huggingface.co/datasets/alexandrainst/m_mmlu, Multilingual ARC at https://huggingface.co/datasets/alexandrainst/m_arc, and Multilingual Hellaswag at https://huggingface.co/datasets/alexandrainst/m_hellaswag.
  • HumanEval+ and MBPP+ (liu2023your) automatically augment the test-cases of the original HumanEval (chen2021evaluating) and MBPP (austin2021program) datasets for code generation assessment. These benchmarks evaluate models’ ability to synthesize programs from docstrings and natural language descriptions, respectively. HumanEval+ and MBPP+ provide 8080x/3535x more tests than the originals, with test splits consisting of $164$ and $378$ programming tasks, respectively. These augmented datasets are publicly provided as follows: HumanEval+ at https://huggingface.co/datasets/evalplus/humanevalplus and MBPP+ at https://huggingface.co/datasets/evalplus/mbppplus.
  • WildGuardTest (han2024wildguard) is a large-scale and carefully balanced multi-task safety moderation dataset. The dataset contains $1,725$ harmful and unharmful samples covering vanilla (direct) and adversarial prompts. However, we only evaluate the merge models’ ability to detect harm using $754$ harmful samples. The dataset is publicly available at https://huggingface.co/datasets/allenai/wildguardmix.
  • HarmBench (mazeika2024harmbench) is a standardized evaluation framework for automated red teaming methods. HarmBench contains $400$ textual behaviors, split into $320$ behaviors for test and $80$ behaviors for validation. These behaviors are designed to violate laws or norms, such that LLMs should not exhibit them. Each behavior is further specified with two types of categorization: semantic and functional categories. Semantic category describes the type of harmful behavior, including cybercrime, copyright violations, and generating misinformation. Functional category describes properties of behaviors, which help measure LLM’s robustness. The dataset is publicly available at https://github.com/nouhadziri/safety-eval-fork/blob/main/evaluation/tasks/generation/harmbench/harmbench_behaviors_text_test.csv.
  • DoAnythingNow (shen2024anything) contains jailbreak prompts (spanning from December 2022 to December 2023), which are exploited by malicious users to bypass the safeguards and elicit harmful content from LLMs. These prompts are collected from four prominent platforms commonly used for prompt sharing: Reddit, Discord, websites, and open-source datasets. Following he2025mergebench, we evaluate the merge models on a subset of $300$ jailbreak prompts created from February 2023 to April 2023. The dataset is publicly available at https://github.com/nouhadziri/safety-eval-fork/blob/main/evaluation/tasks/generation/do_anything_now/do_anything_now_jailbreak.jsonl.
  • XSTest (rottger2024xstest) evaluates whether models’ safeguards are exaggerated. XSTest is inspired by a circumstance where models often struggle to balance helpfulness and harmlessness: a clearly safe prompt is even refused if it uses similar language to unsafe ones or mentions sensitive topics. XSTest contains $450$ test prompts: $250$ safe prompts and $200$ unsafe prompts. The dataset is publicly available at https://github.com/nouhadziri/safety-eval-fork/blob/main/evaluation/tasks/generation/xstest/exaggerated_safety.json.
B.2 Models
Vision classification models.
@@ B.2 Models
We employ off-the-shelf fine-tuned checkpoints from the previous work (tang2024fusionbench), covering three architectures of pre-trained CLIP model (radford2021learning): ViT-B/32, ViT-B/16, and ViT-L/14. The indicators $32$, $16$, and $14$ mean the size of patches in pixels that an input image is divided into. We only merge the vision encoding part of these architectures, while the text encoding part is kept unchanged. The number of parameters for the vision encoding part of these architectures is 87.587.5M, 85.885.8M, and 303303M, respectively. Fine-tuned checkpoints on the standard 8-task benchmark are publicly provided as follows: ViT-B/32 at https://huggingface.co/collections/tanganke/clip-vit-b-32-on-the-eight-image-classication-tasks, ViT-B/16 at https://huggingface.co/collections/tanganke/clip-vit-b-16-on-the-eight-image-classification-tasks, and ViT-L/14 at https://huggingface.co/collections/tanganke/clip-vit-l-14-on-the-eight-image-classification-tasks.
Language generation models.
We employ off-the-shelf fine-tuned checkpoints from the previous work (he2025mergebench), covering two Llama 3 variants (grattafiori2024llama3herdmodels): Llama-3.2-3B and Llama-3.1-8B. For further details on these checkpoints, we refer readers to Section 3.3 of he2025mergebench. Fine-tuned checkpoints are publicly provided as follows: Llama-3.2-3B at https://huggingface.co/collections/MergeBench/llama-32-3b-models and Llama-3.1-8B at https://huggingface.co/collections/MergeBench/llama-31-8b-models.
B.3 Merging Methods
@@
    • Model Soups (wortsman2022modelsoupsaveragingweights) is the most straightforward approach that simply takes the average of candidate models’ parameters to produce a merge model.
    • Task Arithmetic (ilharco2023editingmodelstaskarithmetic) introduces a concept called “task vector”, which is the difference between the fine-tuned model’s parameters and the pre-trained parameters. A multi-tasking task vector is defined as the sum of those task vectors and is scaled by a coefficient before being added back to the pre-trained model’s parameters to produce a merge model.
    • TIES-Merging (yadav2023tiesmergingresolvinginterferencemerging) proposes to trim the small values of task vectors, then resolve sign conflicts before adding back to the pre-trained parameters.
    • TSV-M (gargiulo2025task) compresses the task vectors using singular value decomposition (SVD) to reduce the interference between task vectors at the layer level before merging.
    • DOGE TA (wei2025modeling) is a variant of Task Arithmetic, where DOGE, an iterative algorithm minimizing the gap between the merge model and the candidates while retaining the shared knowledge, is integrated.
    • Iso-C and Iso-CTS (marczakno). The former applied SVD on the merge task vector to identify the directions amplified by multiple tasks, i.e., common subspace. The latter further incorporates task-specific subspaces for retaining unique task features.
@@
    • Fisher Merging (matena2022mergingmodelsfisherweightedaveraging) produces the merge models by taking the weighted average of candidate models, with the weighting factors determined by the Fisher information matrices.
    • RegMean (jin2025datalessknowledgefusionmerging) proposes a closed-form solution for merging multiple linear layers, then applies this idea to the transformer models.
@@
    • Layer-wise AdaMerging (yang2024adamergingadaptivemodelmerging) adaptively learns the merging coefficients introduced by Task Arithmetic in the layer-wise or task-wise manner by using unsupervised entropy minimization on unlabeled test datasets.
    • DOGE AM (wei2025modeling) is another variant of AdaMerging, where DOGE is integrated.
Appendix C Experimental Details
C.1 Normalized Accuracy Metric
@@
C.2 Language Tasks’ Metrics
We provide the task-specific metrics for merging evaluation on language tasks in Table 7.
Domain Dataset Metric
Instruction following IFEval (zhou2023instruction) Prompt level accuracy
Mathematics GSM8K (cobbe2021training) Exact match (8-shot Chain-of-Thought)
Multilingual understanding
(on French, Spanish, German, and Russian)
Multilingual MMLU (lai-etal-2023-okapi) Accuracy
Multilingual ARC (lai-etal-2023-okapi) Normalized accuracy
Multilingual Hellaswag (lai-etal-2023-okapi) Normalized accuracy
Coding HumanEval+ (liu2023your) Pass@1
MBPP+ (liu2023your) Pass@1
Safety WildGuardTest (han2024wildguard) Refuse to answer
HarmBench (mazeika2024harmbench) Refuse to answer
DoAnythingNow (shen2024anything) Refuse to answer
XSTest (rottger2024xstest) Accuracy
Table 7: Task-specific metrics for merging evaluation on language tasks.
C.3 Hyperparameters
Method $0.10$ $0.15$ $0.20$ $0.25$ $0.30$ $0.35$ $0.40$ $0.45$ $0.50$ $0.55$ $0.60$ $0.65$ $0.70$ $0.75$ $0.80$ $0.85$ $0.90$ $0.95$ $1.00$
ViT-B/32
RegMean $76.7$ $78.4$ $79.7$ $80.7$ $81.1$ $81.9$ $82.7$ $83.3$ $83.7$ $84.2$ $84.6$ $84.9$ $85.8$ $85.9$ $86.6$ $86.9$ $87.6$ $87.6$ $3.8$
RegMean++ $78.5$ $80.0$ $81.4$ $82.1$ $83.2$ $83.6$ $84.5$ $85.0$ $85.8$ $86.2$ $86.9$ $87.4$ $87.8$ $88.0$ $88.9$ $89.4$ $89.8$ $90.1$ $5.0$
ViT-B/16
RegMean $80.7$ $82.0$ $82.9$ $83.6$ $84.1$ $84.9$ $85.7$ $85.8$ $86.6$ $87.0$ $87.5$ $88.0$ $88.4$ $88.6$ $89.0$ $89.1$ $89.6$ $90.3$ $4.1$
RegMean++ $82.3$ $83.8$ $84.9$ $85.6$ $86.2$ $86.9$ $87.2$ $87.8$ $88.1$ $88.5$ $88.9$ $89.4$ $90.0$ $90.2$ $90.7$ $91.0$ $91.3$ $91.6$ $4.7$
ViT-L/14
RegMean $87.1$ $88.0$ $88.7$ $89.4$ $89.7$ $90.2$ $90.6$ $91.1$ $91.5$ $91.7$ $92.2$ $92.4$ $92.5$ $92.8$ $93.1$ $93.4$ $93.7$ $94.0$ $4.8$
RegMean++ $88.0$ $89.2$ $89.8$ $90.6$ $91.0$ $91.6$ $91.9$ $92.2$ $92.5$ $92.6$ $93.0$ $93.3$ $93.5$ $93.7$ $94.0$ $94.5$ $94.7$ $94.8$ $4.6$
Table 8: Performance of 88-task scenario merging on held-out validation sets when varying the scaling factor $\alpha$ for non-diagonal items of the inner-product matrices.
@@ C.3 Hyperparameters
Both require two hyperparameters: the number of samples per task and the scaling factor $\alpha$. As illustrated in main text Figure 7 and Figure 8, increasing the number of samples for each task gives a relatively small improvement on performance across architectures. We choose to use $256$ samples for calculating the task’s inner-product matrices with a batch size of $32$ as default. For the scaling factor $\alpha$, Table 8 shows that a higher $\alpha$ delivers better performance. However, when $\alpha=1.0$, i.e., no scaling applied, the degradation happens and the overall accuracy is almost zeroed out. This phenomenon is consistent with the insight from jin2025datalessknowledgefusionmerging. Therefore, $\alpha=0.95$ is the optimal value.
@@
  • For Task Arithmetic (ilharco2023editingmodelstaskarithmetic), the merging coefficient $\lambda=0.3$ is used.
  • For TIES-Merging (yadav2023tiesmergingresolvinginterferencemerging), top-$20\%$ highest-magnitude parameters are retained for each task vector, then these trimmed task vectors are merged with the merging coefficient $\lambda=0.3$.
  • For TSV-M (gargiulo2025task), the task scaling factor $\alpha=1.0$ is used.
  • For DOGE TA (wei2025modeling), the modification vector $\Delta$ is optimized on $400$ iterations with a learning rate 1e41e-4 via Adam optimizer  (kingma2014adam), the global magnitude of merging coefficient $\eta=0.07$, the shared subspace basis size is set as the rank of the shared subspace divided by $6$, and top-$30\%$ highest-magnitude parameters are retained for each task vector.
  • For Iso-C (marczakno), the task scaling factor $\alpha$ is set to $1.30$, $1.40$, and $1.50$ for ViT-B/32, ViT-B/16, and ViT-L/14, respectively.
  • For Iso-CTS (marczakno), the task scaling factor $\alpha$ is set to $1.50$, $1.60$, and $1.90$ for ViT-B/32, ViT-B/16, and ViT-L/14, respectively; the size of the common subspace is set as its rank multiplied by $0.8$.
  • For Fisher Merging (matena2022mergingmodelsfisherweightedaveraging), the number of samples per task is $256$ and the batch size is $32$ for calculating the Fisher information matrix.
  • For Layer-wise AdaMerging (yang2024adamergingadaptivemodelmerging), the Adam optimizer is used with a learning rate of 1e31e-3 for updating the merging coefficients on $1,000$ iterations with the batch size of $16$.
  • For DOGE AM (wei2025modeling), its hyperparameters are the same as DOGE TA and Layer-wise AdaMerging.
C.4 Details for Experimental Settings
Order of tasks for sustainability evaluation.